OpenStudy (anonymous):

(multiplicative inverse,multiplicative identity, multiplicative property of zero, additive inverse, additive identity) 4*(1/4) = 1 3+0 = 3 10*0 = 0 5 + (-5) = 0 8*1 = 8

OpenStudy (anonymous):

4*(1/4) = 1 additive inverse 3+0 = 3 additive identity 10*0 = 0 multiplicative property of zero 5 + (-5) = 0 8*1 = 8 multiplicative inverse

OpenStudy (anonymous):

4*(1/4) = 1 ant that multiplicative identity

OpenStudy (skullpatrol):

yes 4*(1/4) is the multiplicative inverse

OpenStudy (anonymous):

is it inverse or identity

OpenStudy (triciaal):

inverse

OpenStudy (skullpatrol):

the identities involve the identity elements of 0 and 1

OpenStudy (anonymous):

4*(1/4) = 1 multiplicative inverse 3+0 = 3 additive identity 10*0 = 0 multiplicative property of zero 5 + (-5) = 0 additive inverse 8*1 = 8 multiplicative inverse multiplicative identity

OpenStudy (skullpatrol):

when 0 is added to any number the sum is identical to that number

OpenStudy (skullpatrol):

when 1 is multiplied by any number the product is identical to that number

OpenStudy (skullpatrol):

4*(1/4) = 1 multiplicative inverses 3+0 = 3 additive identity 10*0 = 0 multiplicative property of zero 5 + (-5) = 0 additive inverses 8*1 = 8 multiplicative identity

OpenStudy (skullpatrol):